### v1

# Deep Modularity Networks with Diversity–Preserving Regularization

Abstract
Graph clustering plays a crucial role in graph representation learning but often faces challenges in achieving feature-space diversity. While Deep Modularity Networks (DMoN) leverage modularity maximization and collapse regularization to ensure structural separation, they do not explicitly encourage diversity in the feature space among clusters. We address this limitation by proposing Deep Modularity Networks with Diversity-Preserving Regularization (DMoN–DPR), which introduces three novel regularization terms: distance-based for inter-cluster separation, variance-based for intra-cluster diversity, and entropy-based for balanced assignments. Our method enhances clustering performance on benchmark datasets, namely Cora, CiteSeer, PubMed, Coauthor CS, and Coauthor Physics, achieving significant improvements in Normalized Mutual Information (NMI), and F1 scores. These results demonstrate the effectiveness of incorporating diversity-preserving regularizations in creating meaningful and interpretable clusters, especially in feature-rich datasets.
1 Introduction
Graph clustering is a crucial problem within the graph representation learning domain, and is essential for various applications, including but not limited to community detection in social networks (Perozzi et al., 2014; Xiao et al., 2015), data exploration (Perozzi and Akoglu, 2018), and functional module identification in biological networks (Jin et al., 2021). Consequently, there has been a surge in methods aimed at enhancing graph clustering performance. Recent advancements have included techniques like Graph Neural Networks (GNNs) (Scarselli et al., 2008), which leverage node features and graph structure for representation learning, often through unsupervised training (Tsitsulin et al., 2023). In this context, graph pooling methods have also become prominent, as they provide a way to coarsen graphs by aggregating nodes into clusters (Cangea et al., 2018). Yet, some of these methods such as DiffPool (Ying et al., 2018) and MinCutPool (Bianchi et al., 2019) were found to be computationally costly, or too rigid, leading to poor convergence (Tsitsulin et al., 2023).
To address these limitations, Deep Modularity Networks (DMoN) (Tsitsulin et al., 2023) were introduced, combining spectral modularity maximization (Newman, 2006) with collapse regularization (Tsitsulin et al., 2023) to avoid trivial clustering solutions. While DMoN improves structural separation, it does not explicitly enforce diversity among clusters in the feature space. This limitation can result in structurally distinct but feature-wise homogeneous clusters, reducing effectiveness in applications requiring diverse and meaningfully differentiated clusters.
Present Work
To overcome the limitations of DMoN and enhance feature-space diversity in graph clustering, we propose an extension called Deep Modularity Networks with Diversity-Preserving Regularization (DMoN–DPR). Our method augments the original DMoN optimization objective by introducing additional regularization terms specifically designed to promote diversity among clusters. These include a distance-based regularization term, which penalizes clusters with centroids too close in feature space to encourage distinct separation; a variance-based regularization term, which maximizes the variance within each cluster’s assignment probabilities to capture a broader range of node features; and an entropy-based regularization term, which ensures a more balanced and uniform node distribution by maximizing the entropy of cluster assignments.
Our proposed DMoN-DPR method addresses the shortcomings of existing approaches by explicitly enforcing feature-space diversity and ensuring that clusters remain well-separated, all while preventing them from becoming overly homogeneous in their node features. As a result, the clusters produced are both structurally cohesive and feature-wise distinct. Following the evaluation framework in (Tsitsulin et al., 2023), we demonstrate through extensive experiments that our method improves clustering performance on the benchmark datasets, noticeably on Coauthor CS and Coauthor Physics datasets, which benefit from this enriched representation due to their rich feature spaces.
2 Related Work
Early methods for clustering on graphs often treated node features independently from graph structure. A straightforward example is k-means(features) (Lloyd, 1982), which clusters nodes solely based on attributes and disregards their connectivity. While simple, this approach often overlooks crucial structural relationships. To incorporate at least some graph information, k-means(DeepWalk) (Perozzi et al., 2014) combines node attributes with random-walk-based embeddings. This introduces a basic integration of graph structure but still relies on a two-step process of embedding and clustering. A more advanced alternative is k-means(DGI) (Tsitsulin et al., 2023), which employs node representations learned via Deep Graph Infomax (Veličković et al., 2018) and thereby captures both local and global context. However, because these k-means strategies only merge attributes and structure post hoc—rather than jointly optimizing them—they may lag behind end-to-end trainable methods that unify both graph topology and node features in a single framework.
Moving toward more integrated approaches, Defferrard et al. (2016) introduced one of the first spectral graph convolution methods based on Chebyshev polynomials. Their technique approximates the spectral filters necessary for convolutional operations on graphs, reducing the computational cost of graph Fourier transforms and providing a scalable foundation for subsequent GNN designs. This seminal work illustrated how learned transformations on graph-structured data could drastically improve learning efficiency and accuracy, motivating the community to delve deeper into neural approaches.
Beyond simply learning node-level representations, pooling mechanisms aim to capture hierarchical structures within graphs. Neural Overlapping Community Detection (NOCD) (Shchur and Günnemann, 2019) applies a probabilistic model that directly optimizes the likelihood of reconstructing the graph. Although effective in certain scenarios, NOCD may face scalability challenges and underutilizes node features in complex datasets. One of the earliest neural pooling frameworks, DiffPool (Ying et al., 2018), introduced a soft-clustering strategy that learns cluster assignments in an end-to-end manner. Despite its strong performance, its quadratic time complexity renders it less practical for large-scale graphs (Tsitsulin et al., 2023).
To address computational bottlenecks and overcome potential overfitting, MinCut Pooling (Bianchi et al., 2020) reimagines the clustering problem through a spectral lens, incorporating normalized minCUT objectives with orthogonality constraints. While these constraints can improve cluster quality, they frequently lead to convergence issues. A related simplified variant, Ortho (Bianchi et al., 2020), focuses exclusively on orthogonality regularization, thus discarding many graph-specific objectives. This omission can produce suboptimal results, as the method forgoes some of the structural cues integral to good cluster formation. In another line of work, SAGPool (Lee et al., 2019) introduces a self-attention mechanism that ranks node importance, thereby selecting the most informative nodes during pooling. Although SAGPool effectively combines feature and topology considerations, computing attention coefficients can increase training costs, particularly for large graphs.
Orthogonal to pooling-based approaches, Deep Graph Infomax (DGI) (Velickovic et al., 2019) pioneered an approach that maximizes mutual information between local substructures and a global summary of the graph. This method facilitates robust, label-agnostic node embeddings. The framework was later extended by InfoGraph (Sun et al., 2019), which applies a similar principle at the graph-level, yielding state-of-the-art results for unsupervised and semi-supervised tasks by capturing interactions between subgraphs and a holistic representation.
Most recently, Deep Modularity Networks (DMoN) (Tsitsulin et al., 2023) have demonstrated a powerful alternative by directly incorporating modularity optimization with collapse regularization in an end-to-end trainable pipeline. DMoN bridges traditional graph community detection and neural feature extraction, achieving strong empirical performance on both synthetic and real-world graphs. By jointly optimizing for high modularity and balancing cluster assignment, DMoN outperforms many existing approaches in terms of normalized mutual information (NMI) and modularity scores, thus setting a new standard for clustering-oriented GNNs.
3 Deep Modularity Networks with Diversity–Preserving Regularization
DMoN tackles the issues of previous techniques by leveraging an optimization objective that combines insights from spectral modularity maximization (Newman, 2006) with a unique regularization term called collapse regularization (Tsitsulin et al., 2023). Specifically, DMoN encodes the cluster assignments, represented as a soft assignment matrix $C$ by using a softmax function over the output of a GNN, allowing differentiation during optimization. For each node, a soft cluster assignment $C$ is computed as follows:
$C=\text{softmax}(\text{GCN}(\tilde{A},X))$ (1)
where GCN is a multi-layer graph convolutional network, $\tilde{A}$ is the normalized adjacency matrix $\tilde{A}=D^{-\frac{1}{2}}AD^{-\frac{1}{2}}$, and $X$ represents the node features. The objective function of DMoN, $L_{\text{DMoN}}$, combines a modularity term with a collapse regularization term to optimize the clustering. The modularity term measures the quality of the cluster assignments by maximizing the density of intra-cluster edges relative to a null model, while the collapse regularization prevents trivial solutions where all nodes are assigned to the same cluster. The objective function is formulated as:
$L_{\text{DMoN}}(C;A)=\underbrace{-\frac{1}{2m}\text{Tr}(C^{\top}BC)}_{\text{Modularity term}}+\underbrace{\sqrt{\frac{k}{n}}\left\|\sum_{i}C_{i}^{\top}\right\|_{F}-1}_{\text{Collapse Regularization term}}$ (2)
where $\frac{1}{2m}\text{Tr}(C^{\top}BC)$ is the modularity term (measuring how well clusters are internally connected compared to random chance), $B=A-\frac{dd^{\top}}{2m}$ is the modularity matrix, $d$ is the degree vector, $m$ is the total number of edges, $\sqrt{\frac{k}{n}}\left\|\sum_{i}C_{i}^{\top}\right\|_{F}-1$ represents the collapse regularization term, which discourages the formation of trivial clusters by penalizing the Frobenius norm of the assignment matrix $C$, $\|\cdot\|_{F}$ denotes the Frobenius norm, and $k$ is the number of clusters. This regularization term encourages balanced cluster assignments, thereby improving the quality of clustering and avoiding degenerate solutions.
Building upon the DMoN framework, we augment the DMoN objective function $L_{\text{DMoN}}$ with three new regularization terms, resulting in the following enhanced objective function:
$L_{\text{DMoN--DPR}}(C;A)=L_{\text{DMoN}}+W_{\text{dist}}\cdot L_{\text{DPR}}^{\text{Distance}}+W_{\text{var}}\cdot L_{\text{DPR}}^{\text{variance}}+W_{\text{entropy}}\cdot L_{\text{DPR}}^{\text{entropy}}$ (3)
Here, $W_{\text{dist}}$, $W_{\text{entropy}}$, and $W_{\text{var}}$ are hyperparameters that control the influence of each regularization term. The additional terms are designed to promote inter-cluster separation, intra-cluster diversity, and balanced cluster assignments.
Distance-Based Regularization
Inspired by SimCLR (Chen et al., 2020), which showed that contrastive loss promotes well-separated clusters, the distance-based regularization term encourages distinct cluster centroids in feature space. It is defined as:
$L_{\text{DPR}}^{\text{Distance}}=\frac{1}{k(k-1)}\sum_{i=1}^{k}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{k}\text{ReLU}\left(\epsilon-\|\mu_{i}-\mu_{j}\|_{2}^{2}\right)$ (4)
where $k$ is the number of clusters, $\mu_{i}$ and $\mu_{j}$ are the centroids of clusters $i$ and $j$, computed as $\mu_{i}=\frac{\sum_{v=1}^{n}C_{vi}X_{v}}{\sum_{v=1}^{n}C_{vi}}$, where $X_{v}$ is the feature vector of node $v$, $\|\cdot\|_{2}$ denotes the Euclidean norm, $\epsilon$ is a predefined threshold that sets the minimum acceptable distance between cluster centroids, and ReLU is the Rectified Linear Unit function, ensuring that only distances below $\epsilon$ contribute to the loss. By penalizing pairs of clusters whose centroids are closer than $\epsilon$, clusters can be effectively pushed apart in the feature space. By encouraging greater separation between clusters, it enhances inter-cluster diversity and reduces overlap, leading to more distinguishable and meaningful clusters.
Variance-Based Regularization
The variance-based regularization term aims to increase the diversity within each cluster by maximizing the variance of node assignments in each cluster:
$L_{\text{DPR}}^{\text{variance}}=-\frac{1}{k}\sum_{i=1}^{k}\text{Var}(C_{:i})$ (5)
[…]
Entropy-Based Regularization
The entropy-based regularization term promotes a uniform distribution of nodes across clusters, preventing any single cluster from dominating. It is formulated as:
$L_{\text{DPR}}^{\text{entropy}}=-\frac{1}{n}\sum_{v=1}^{n}\sum_{i=1}^{k}C_{vi}\log(C_{vi}+\delta)$ (6)
where $n$ is the number of nodes, $C_{vi}$ is the soft assignment probability of node $v$ to cluster $i$, $\delta$ is a small constant added to prevent logarithm of zero, ensuring numerical stability. By maximizing the entropy of the cluster assignments, this term encourages uncertainty in the assignments, promoting a more balanced distribution of nodes among clusters. This prevents scenarios where one cluster might contain a majority of the nodes, thus enhancing overall diversity and cluster balance.
Therefore, by integrating these diversity-preserving regularization terms into the DMoN objective, we can simultaneously consider graph structure, cluster balance, inter-cluster separation, and intra-cluster diversity.
4 Experiments
Datasets
To evaluate the effectiveness of our proposed regularization objective, we conducted experiments on the widely used Cora, CiteSeer, and PubMed datasets Yang et al. (2016), as well as the Coauthor CS and Coauthor Physics datasets (Shchur et al., 2018). The Cora, CiteSeer, and PubMed datasets are citation networks where nodes represent papers and edges denote citation relationships, with node labels corresponding to the topic of each paper. The Coauthor CS and Coauthor Physics datasets are co-authorship graphs, where nodes represent authors and edges indicate collaborations, with labels reflecting the field of research. The key characteristics of these datasets are summarized in Table 1.
Table 1: Dataset characteristics, where ($|V|$) is the number of vertices , ($|E|$) is the number of edges, $|X|$ is the number of features, and $|Y|$ is the number of cluster labels.
Dataset $|V|$ $|E|$ $|X|$ $|Y|$
Cora 2708 5278 1433 7
CiteSeer 3327 4614 3703 6
PubMed 19717 44325 500 3
Coauthor CS 18333 81894 6805 15
Coauthor Phys 34493 247962 8415 5
Evaluation Protocol
Following the evaluation protocol outlined in (Tsitsulin et al., 2023), we employed the following metrics to assess performance: graph conductance (C), modularity (Q), Normalized Mutual Information (NMI) with ground-truth labels, and the pairwise F1 measure. These metrics collectively capture the structural quality of clusters and their alignment with ground-truth labels.
Baseline
We select DMoN as our sole baseline because it consistently achieves state-of-the-art results across multiple datasets and metrics, as shown in prior work (Tsitsulin et al., 2023). It outperforms a wide range of established methods—such as k-means(features) (Lloyd, 1982), k-means(DeepWalk) (Perozzi et al., 2014), k-means(DGI) (Veličković et al., 2018), NOCD (Shchur and Günnemann, 2019), DiffPool (Ying et al., 2018), MinCut Pooling (Bianchi et al., 2020), and Ortho (Bianchi et al., 2020)—each representing distinct approaches that prioritize node attributes, graph structure, or both. Unlike traditional clustering methods like k-means (which lacks graph structure awareness) or embedding-based techniques (which emphasize representation learning over explicit clustering), DMoN offers an end-to-end differentiable framework that directly optimizes modularity and uses collapse regularization to avoid trivial solutions. Furthermore, it exhibits superior scalability and computational efficiency compared to pooling-based approaches such as DiffPool and MinCutPool, making it particularly suited for the large, feature-rich datasets in our experiments. By focusing on feature-space diversity within DMoN’s modularity-driven clustering paradigm, we ensure that our enhancements remain aligned with state-of-the-art structural considerations, providing a clear and meaningful evaluation of the improvements introduced by our proposed modifications.
Implementation Details
Similar to (Tsitsulin et al., 2023), one layer of GCN (Kipf and Welling, 2016) with 512 neurons was used to create the graph embedding, followed by the DMoN/DMoN–DPR pooling layer, as illustrated in Figure 1. Likewise, the models were trained for 1000 epochs using Adam optimizer with a learning rate of 0.001. The experiments were run on an Apple M2 Max CPU. The code–implemented by extending the DMoN implementation in PyTorch Geometric (Fey and Lenssen, 2019)– will be made publicly available at www.github.com/TBA.
Figure 1: Overview of the DMoN/DMoN–DPR pipeline.
Hyperparameter Tuning
The weighting coefficients $W_{\text{dist}}$, $W_{\text{entropy}}$, and $W_{\text{var}}$, as well as $\epsilon$ are hyperparameters that need to be tuned based on the specific dataset and desired clustering behavior. They control the trade-off between structural modularity and diversity preservation. To optimize these hyperparameters, we conducted evaluations on the Cora, CiteSeer, PubMed, Coauthor CS and Coauthor Physics datasets using 10 random seeds, similar to how it was done in (Tsitsulin et al., 2023). The seeds that resulted in the best performance when using the DMoN pooling layer were selected. For each dataset, we tuned $\epsilon$ (for the distance weight $W_{\text{dist}}$), $W_{\text{var}}$ (variance weight), and $W_{\text{entropy}}$ (entropy weight) independently. To find the best $\epsilon$, the variance and entropy weights were set to zero, and $\epsilon$ was varied from 10 to $10^{-5}$. Similarly, $W_{\text{var}}$ and $W_{\text{entropy}}$ were tuned by fixing the other weights to zero and varying them across {1, 0.1, 0.01, 0.001}. The best weights identified were used to construct various DMoN-DPR models, including DMoN (baseline), DMoN–DPR(D) (distance), DMoN–DPR(V) (variance), DMoN–DPR(E) (entropy), and combinations: DMoN–DPR(DV) (distance and variance), DMoN–DPR(DE) (distance and entropy), DMoN–DPR(VE) (variance and entropy), and DMoN–DPR(DVE) (distance, variance, and entropy). This systematic approach allowed us to assess the impact of each regularization term and their combinations on model performance. Ablation studies are presented in Section A.
5 Results
Table 2: Comparison of clustering methods on three datasets (Cora, Citeseer, Pubmed). Values are in percentage.
Method Cora Citeseer Pubmed
Graph Labels Graph Labels Graph Labels
$C\downarrow$ $Q\uparrow$ NMI $\uparrow$ F1 $\uparrow$ $C\downarrow$ $Q\uparrow$ NMI $\uparrow$ F1 $\uparrow$ $C\downarrow$ $Q\uparrow$ NMI $\uparrow$ F1 $\uparrow$
DMoN 11.00 72.39 43.65 46.24 4.92 75.72 30.26 42.17 8.66 57.03 22.68 45.18
DMoN–DPR(D) 10.23 72.97 45.31 50.00 4.75 75.64 30.61 42.76 8.66 57.02 22.68 45.19
DMoN–DPR(V) 11.64 70.80 44.27 48.18 4.93 75.59 30.46 42.26 8.66 57.02 22.68 45.18
DMoN–DPR(E) 11.00 72.37 43.63 46.25 5.02 75.43 30.52 42.54 8.67 57.02 22.68 45.19
DMoN–DPR(DV) 10.37 72.68 45.41 50.13 4.76 75.51 30.59 42.85 8.66 57.02 22.68 45.19
DMoN–DPR(DE) 10.33 72.88 45.18 49.88 4.86 75.35 30.77 43.27 8.66 57.02 22.67 45.19
DMoN–DPR(VE) 11.64 70.77 44.15 48.09 5.09 75.26 30.86 42.86 8.66 57.01 22.68 45.18
DMoN–DPR(DVE) 10.41 72.65 45.43 50.06 4.85 75.14 30.77 43.42 8.66 57.02 22.67 45.18
Table 3: Comparison of clustering methods on two datasets (Coauthor CS, Coauthor Physics). Values are in percentage.
Method Coauthor CS Coauthor Phys
Graph Labels Graph Labels
$C\downarrow$ $Q\uparrow$ NMI $\uparrow$ F1 $\uparrow$ $C\downarrow$ $Q\uparrow$ NMI $\uparrow$ F1 $\uparrow$
DMoN 18.80 72.43 69.45 58.83 13.81 63.35 52.40 46.80
DMoN–DPR(D) 19.01 72.17 70.56 61.09 16.45 56.93 53.99 50.96
DMoN–DPR(V) 18.89 72.41 69.76 59.10 13.45 59.38 54.03 51.89
DMoN–DPR(E) 19.73 70.97 71.34 60.40 13.80 60.04 53.17 50.05
DMoN–DPR(DV) 18.95 72.26 71.33 61.81 13.46 56.85 54.09 56.36
DMoN–DPR(DE) 19.88 70.56 70.84 61.76 14.22 56.10 54.71 56.94
DMoN–DPR(VE) 19.70 71.01 71.45 60.56 12.85 55.72 50.14 54.52
DMoN–DPR(DVE) 19.74 70.68 71.08 62.41 13.46 55.27 54.31 58.86
5.1 Effectiveness of DMoN–DPR in Clustering
To assess the effectiveness of DMoN–DPR, we compare the mean values of conductance ($C$), modularity ($Q$), NMI, and pairwise F1 measure across 10 seeds achieved with DMoN–DPR to those obtained with DMoN, as done in (Tsitsulin et al., 2023) (full results can be found in Section B). Tables 2 (Cora, CiteSeer, PubMed) and 3 (Coauthor CS, Coauthor Physics) summarize the results, where graph metrics (C, Q) do not rely on ground-truth labels, whereas ‘Labels’ metrics (NMI, F1) compare to known class labels.
Cora, CiteSeer, and PubMed.
On Cora, where the diversity of node features and class labels is relatively high, incorporating the distance-based term (D) consistently enhances performance over the baseline DMoN, especially on label-based metrics (NMI and F1). The combinations that include the distance term—such as DMoN–DPR(DV), DMoN–DPR(DE), and DMoN–DPR(DVE)—not only improve cluster separability but also boost label-based performance, with DMoN–DPR(DV) and DMoN–DPR(DVE) achieving some of the best F1 and NMI scores. The entropy (E) and variance (V) terms alone are less impactful, but when paired with distance, they synergize to produce more balanced and diverse clusters. On CiteSeer, which also has a reasonably complex feature space, the introduction of these regularizations yields moderate gains, with combinations that include the entropy term slightly improving label-based metrics. In contrast, PubMed, which has a less diverse feature space, sees only marginal benefits. Here, the baseline DMoN already performs well, and the added terms provide limited additional advantage.
Coauthor CS and Coauthor Physics.
We observe a similar pattern on the two Coauthor datasets. Both of these datasets have richer feature spaces than Cora, CiteSeer, and PubMed; however, Coauthor Physics stands out with the largest number of features among all the datasets under consideration (8,415). As shown in Table 3, this higher dimensionality helps amplify the benefits of our diversity-preserving regularizations, especially those involving the distance-based term (D). Notably, on Coauthor Physics, the pairwise F1 score jumps from 46.80 (DMoN) to 58.86 (DMoN–DPR(DVE)), indicating that combining distance, variance, and entropy can substantially enhance the separability of clusters and the alignment with true labels.
On Coauthor CS, our proposed regularizations also offer improvements over baseline DMoN, most evident when distance (D) is paired with either variance (V) or entropy (E). For example, F1 increases from 58.83 (DMoN) to 62.41 (DMoN–DPR(DVE)). Although these gains are meaningful, they are somewhat more moderate than those on Coauthor Physics—likely due to differing distributions of class labels and feature dimensions. Overall, these findings suggest that introducing distance-, variance-, and entropy-based regularizations becomes increasingly beneficial as the dimensionality and diversity of node features grow, leading to more pronounced performance boosts on datasets such as Coauthor Physics.
Analyzing the Influence of Regularization on Clustering Performance.
Our experiments show that the impact of the distance-based, variance-based, and entropy-based regularizations depends heavily on both the feature richness and label diversity of each dataset. On Coauthor CS, and Coauthor Physics, which have relatively higher-dimensional or more varied feature spaces, we observe notable gains in NMI and F1 upon adding these diversity-promoting terms—especially the distance-based (D) regularization, which strongly boosts inter-cluster separation. For instance, DMoN–DPR(D) and its variants (e.g., DMoN–DPR(DV), DMoN–DPR(DE), DMoN–DPR(DVE)) yield significant improvements in label alignment on Coauthor Physics, where the F1 score rises by over 10 percentage points compared to baseline DMoN.
Meanwhile, the variance-based (V) term is helpful in ensuring clusters capture a wider range of node features, an effect that is especially apparent on datasets like Coauthor Physics. Although the single V-only variant often yields moderate gains on datasets such as Cora, it generally performs best when combined with the distance or entropy regularizers. The entropy-based (E) term aims to produce more balanced cluster distributions, proving modestly beneficial on datasets like CiteSeer and effective on Coauthor CS and Coauthor Physics; however, on PubMed, which has lower feature dimensionality, we see little to no improvement (the numbers remain nearly unchanged from DMoN). This indicates that entropy-based balancing may not substantially help in more homogeneous datasets.
Moreover, combining all three terms (DMoN–DPR(DVE)) often yields the highest F1 scores across multiple datasets (e.g., a leap from 58.83 to 62.41 on Coauthor CS and 46.80 to 58.86 on Coauthor Physics). These results suggest that multiple regularizers can work in synergy, but the exact benefits depend on a dataset’s inherent complexity. Overall, the findings highlight the importance of tailoring regularization strategies to both the graph structure and feature diversity of the data to achieve the most effective clustering outcomes.
Trade-Offs Between Structural and Label-Based Metrics
Our results highlight a trade-off between metrics that prioritize graph structure (e.g., Modularity ($Q$) and Conductance ($C$)) and those that evaluate label alignment (e.g., NMI and F1). This trade-off emerges because diversity-preserving regularization terms influence both the feature and label spaces of clusters, leading to distinct outcomes depending on the dataset characteristics and application needs.
For datasets with rich feature spaces (e.g., Coauthor CS and Coauthor Physics), the diversity-enhancing terms significantly boost label-based metrics like NMI and F1. These gains stem from the formation of more homogeneous clusters that align closely with ground-truth labels. However, this improvement in label alignment can come at the expense of structural metrics such as modularity, which reward densely connected clusters in the graph topology. For instance, while DMoN–DPR(DVE) achieves the highest F1 scores on Coauthor Physics, it exhibits a decrease in modularity compared to baseline DMoN.
Conversely, on datasets with less diverse or lower-dimensional features (e.g., PubMed), the structural metrics are less affected by the introduction of diversity-preserving regularization terms, and the benefits to label-based metrics are more modest. In such cases, the original DMoN, which focuses on structural cohesion, remains a competitive choice for purely unsupervised clustering tasks.
Ultimately, the decision to prioritize label alignment or structural metrics should be informed by the application context. For tasks involving semi-supervised learning or node classification, where accurate label recovery is essential, DMoN–DPR offers clear advantages. In contrast, for community detection or unsupervised settings where graph structure is paramount, the baseline DMoN may be preferable. This balance between competing objectives underscores the need to tailor clustering approaches to the specific demands of the dataset and task at hand.
Figure 2: Visualization of clusters (circles) vs. ground truth labels (crosses) formed by DMoN on (a) Cora, (b) CiteSeer, (c) PubMed, (d) Coauthor CS, and (e) Coauthor Physics, and by DMoN–DPR on (f) Cora, (g) CiteSeer, (h) PubMed, (i) Coauthor CS, and (j) Coauthor Physics datasets.
5.2 Visualization of Clusters
The clustering results visualized in Figure 1 demonstrate that DMoN–DPR generally outperforms DMoN across the benchmark datasets Cora, CiteSeer, PubMed, Coauthor CS, and Coauthor Physics. For Cora, DMoN–DPR produces tighter, more well-separated clusters with improved alignment to the ground truth labels, contrasting with the more scattered and less defined clusters formed by DMoN. Similarly, for CiteSeer, DMoN–DPR exhibits better cluster compactness and reduced overlap, though the gains are somewhat less pronounced than in Cora. On the other hand, PubMed remains challenging for both methods, resulting in notable overlap and scattered clusters; yet, DMoN–DPR still provides slightly tighter groupings. Across these three datasets, the diversity-preserving regularization in DMoN–DPR effectively enhances both inter-cluster distinctiveness and intra-cluster variability, producing more meaningful and interpretable partitions. Turning to the Coauthor CS and Coauthor Physics datasets, a similar trend emerges. While DMoN adequately captures overarching modular structures, it can occasionally struggle to separate nuanced subgroups. In contrast, DMoN–DPR consistently yields more cohesive and better-separated clusters, offering clearer decision boundaries and higher alignment with the true labels. This improvement is particularly notable in Coauthor CS, where DMoN–DPR visibly reduces overlap and sharpens cluster delineations. Although the inherently complex Coauthor Physics dataset can induce some overlap in both methods, DMoN–DPR still produces more robust clusters overall. These findings underscore the effectiveness of promoting diversity within hidden representations, as DMoN–DPR not only excels in forming compact and accurate clusters for datasets with greater feature variability but also remains competitive on more challenging datasets.
5.3 Execution Times and Runtime Analysis
Table 4 reports the execution times (in seconds) of DMoN and DMoN–DPR on five commonly used graph datasets. All experiments were conducted on a CPU to ensure consistent runtime measurement without GPU scheduling effects.
Table 4: Execution times (in seconds) on CPU for DMoN and DMoN–DPR.
Dataset DMoN DMoN–DPR
Cora 32 35
CiteSeer 76 71
PubMed 310 300
Coauthor CS 722 751
Coauthor Physics 2107 1946
Although DMoN–DPR includes additional regularization terms for Distance-based, Entropy-based, and Variance-based constraints, its runtime is sometimes comparable to or even marginally lower than plain DMoN. The reason is that the most time-consuming part of both methods is dominated by adjacency-based operations (e.g., matrix multiplications with the graph Laplacian or adjacency matrix), which scale on the order of $\mathcal{O}(|E|\,k)$ or $\mathcal{O}(N^{2}k)$ for large, dense graphs (where $N$ is the number of nodes, $E$ is the set of edges, and $k$ is the number of clusters). In contrast, the extra DPR calculations—computing entropy across nodes, the variance of cluster assignments, or centroid distances—represent only $\mathcal{O}(N\,k+k^{2}F)$ overhead (with $F$ being the feature dimension) and are typically negligible next to the larger matrix multiplications.
Furthermore, running on CPU can introduce variability from scheduling, caching, and background processes, so small differences in the recorded times may not be statistically meaningful. Thus, although DMoN–DPR performs slightly more computations, the extra cost is overshadowed by the main bottleneck operations, resulting in comparable or even lower measured runtimes in practice.
5.4 Limitations and Future Directions
Despite the improvements introduced by our diversity–preserving regularization, DMoN–DPR has a few notable limitations. First, tuning the regularization weights for distance, variance, and entropy remains a manual process. Determining the best values for these hyperparameters can be time-consuming and sensitive to each dataset’s characteristics. An interesting direction for future work is to develop an automated or adaptive scheme that learns these weights during training, making the method more robust and less reliant on exhaustive search.
Second, while DMoN–DPR yields substantial gains on datasets with a large or diverse feature space, the improvements are often marginal for datasets with lower feature dimensionality (e.g., PubMed). Nonetheless, our visualization in Figure 2 suggests that even in these cases, DMoN–DPR tends to produce slightly more coherent cluster structures compared to the baseline, reflecting the benefit of promoting diversity even when numerical metrics (e.g., NMI, F1) show limited change. Future research could investigate whether alternative or more specialized regularizers can drive greater numerical improvements on lower-dimensional datasets.
6 Conclusion
In this work, we presented DMoN–DPR, an enhanced version of Deep Modularity Networks (DMoN) that incorporates diversity-preserving regularizations to enrich the clustering objective. By introducing distance-, entropy-, and variance-based penalties alongside the original modularity and collapse regularization terms, our approach promotes both inter- and intra-cluster diversity. Our empirical results on benchmark datasets indicate that DMoN–DPR significantly improves alignment with ground-truth labels, particularly in feature-rich datasets. While structural metrics such as modularity and conductance remained competitive, the additional diversity constraints contributed to more interpretable and semantically meaningful cluster formations. Overall, DMoN–DPR achieves a balance between interpretability and structural integrity, making it a powerful solution for graph clustering tasks that demand both topological cohesion and semantic differentiation, marking a significant advancement in feature-aware graph clustering.
Appendix A Ablation Study
A.1 Varying the Epsilon
Figure 3 depicts the ablation study done to find the best value of $\epsilon$ associated with $W_{\text{dist}}$ by setting $W_{\text{var}}$ and $W_{\text{entropy}}$ weights to 0. For the Cora, CiteSeer, PubMed, Coauthor CS, and Coauthor Physics datasets, the best $\epsilon$ values were found to be $0.0001$, $0.0001$, $0.001$, $1.0$, and $10.0$ respectively.
Figure 3: Finding the best value of epsilon ($\epsilon$) by setting $W_{\text{dist}}$ to 1 and $W_{\text{var}}$ and $W_{\text{entropy}}$ to zero on the (a) Cora, (b) CiteSeer, (c) PubMed, (d) Coauthor CS, and (e) Coauthor Physics datasets.
A.2 Varying the Variance Weight
Figure 4 depicts the ablation study done to find the best value of $W_{\text{var}}$ by setting $W_{\text{dist}}$ and $W_{\text{entropy}}$ weights to 0. For the Cora, CiteSeer, PubMed, Coauthor CS, and Coauthor Physics datasets, the best $W_{\text{var}}$ values were found to be $1$/$0.1$, $0.1$, $0.001$, $0.1$, and $1.0$ respectively.
Figure 4: Finding the best value of $W_{\text{var}}$ by setting $W_{\text{dist}}$ and $W_{\text{entropy}}$ to zero on the (a) Cora, (b) CiteSeer, (c) PubMed, (d) Coauthor CS, and (e) Coauthor Physics datasets.
A.3 Varying the Entropy Weight
Figure 5 depicts the ablation study done to find the best value of $W_{\text{entropy}}$ by setting $W_{\text{dist}}$ and $W_{\text{var}}$ weights to 0. For the Cora, CiteSeer, PubMed, Coauthor CS, and Coauthor Physics datasets, the best $W_{\text{entropy}}$ values were found to be $0.001$, $0.01$, $0.001$, $0.1$ and $0.1$ respectively.
Figure 5: Finding the best value of $W_{\text{entropy}}$ by setting $W_{\text{dist}}$ and $W_{\text{var}}$ to zero on the (a) Cora, (b) CiteSeer, (c) PubMed, (d) Coauthor CS, and (e) Coauthor Physics datasets.
Appendix B Full Results
The tables below are indicative of the value of each evaluation metric obtained using DMoN/DMoN–DPR pooling for each individual seed for different datasets. The numbers are all in percentage (%).
B.1 Cora
The following tables list the performance of DMoN and different DMoN–DPR configurations across different seeds on the Cora dataset.
Table 5: Results obtained by using DMoN pooling on the Cora dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN 10.25 73.93 48.71 55.38
550 DMoN 9.95 71.77 40.62 42.17
243 DMoN 10.06 73.74 47.62 53.20
16 DMoN 12.13 70.62 44.77 46.61
716 DMoN 11.92 70.40 37.87 39.65
383 DMoN 11.20 71.80 44.40 52.46
277 DMoN 11.03 72.50 37.07 36.07
274 DMoN 10.12 74.14 45.30 44.74
188 DMoN 11.48 73.36 46.79 49.47
796 DMoN 11.86 71.67 43.30 42.69
Mean 11.00 72.39 43.65 46.24
Variance 0.01 0.02 0.16 0.40
Table 6: Results obtained by using DMoN–DPR pooling, using the best epsilon value of 0.0001, and distance weight of 1 on the Cora dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (D) 10.50 73.78 51.01 57.45
550 DMoN–DPR (D) 10.06 72.20 41.49 44.43
243 DMoN–DPR (D) 9.49 73.58 47.80 54.94
16 DMoN–DPR (D) 10.55 72.20 46.22 49.19
716 DMoN–DPR (D) 10.38 71.10 39.29 43.43
383 DMoN–DPR (D) 9.49 74.11 48.02 54.77
277 DMoN–DPR (D) 10.65 72.45 40.12 42.77
274 DMoN–DPR (D) 9.47 74.41 47.64 52.47
188 DMoN–DPR (D) 10.44 73.75 48.41 54.78
796 DMoN–DPR (D) 11.29 72.13 43.05 45.75
Mean 10.23 72.97 45.31 50.00
Variance 0.00 0.01 0.16 0.31
Table 7: Results obtained by using DMoN–DPR pooling, using the best variance weight of 1 on the Cora dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR 10.99 72.41 51.86 58.92
550 DMoN–DPR (V) 11.10 70.31 40.04 43.09
243 DMoN–DPR (V) 11.75 71.02 47.47 54.54
16 DMoN–DPR (V) 10.99 70.18 45.82 48.81
716 DMoN–DPR (V) 11.08 69.84 40.82 41.88
383 DMoN–DPR (V) 12.13 69.37 42.89 51.12
277 DMoN–DPR (V) 11.42 70.91 37.43 37.64
274 DMoN–DPR (V) 11.05 72.61 46.97 48.40
188 DMoN–DPR (V) 13.11 70.74 45.71 52.26
796 DMoN–DPR (V) 12.79 70.56 43.70 45.13
Mean 11.64 70.80 44.27 48.18
Variance 0.01 0.01 0.18 0.41
Table 8: Results obtained by using DMoN–DPR pooling, using the best entropy weight of 0.001 on the Cora dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (E) 10.25 73.81 49.10 55.86
550 DMoN–DPR (E) 10.06 71.70 40.49 42.06
243 DMoN–DPR (E) 9.91 73.82 48.60 53.72
16 DMoN–DPR (E) 12.32 70.46 44.95 46.88
716 DMoN–DPR (E) 11.86 70.43 37.72 39.66
383 DMoN–DPR (E) 11.01 71.93 44.49 52.83
277 DMoN–DPR (E) 11.16 72.37 37.38 35.98
274 DMoN–DPR (E) 10.17 74.20 44.13 44.02
188 DMoN–DPR (E) 11.41 73.37 46.28 48.86
796 DMoN–DPR (E) 11.88 71.65 43.17 42.63
Mean 11.00 72.37 43.63 46.25
Variance 0.01 0.02 0.16 0.43
Table 9: Results obtained by using DMoN–DPR pooling, using the best epsilon value of 0.0001, distance weight of 1, and variance weight of 0.1 on the Cora dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DV) 10.25 73.42 51.71 58.37
550 DMoN–DPR (DV) 10.17 72.08 41.28 44.08
243 DMoN–DPR (DV) 9.66 73.18 48.08 55.94
16 DMoN–DPR (DV) 10.72 71.69 46.80 49.51
716 DMoN–DPR (DV) 10.52 70.97 39.46 43.32
383 DMoN–DPR (DV) 9.87 73.82 48.11 54.87
277 DMoN–DPR (DV) 10.95 71.90 39.01 42.43
274 DMoN–DPR (DV) 9.61 74.13 48.04 53.26
188 DMoN–DPR (DV) 10.57 73.51 48.47 54.55
796 DMoN–DPR (DV) 11.39 72.06 43.12 44.93
Mean 10.37 72.68 45.41 50.13
Variance 0.00 0.01 0.19 0.36
Table 10: Results obtained by using DMoN–DPR pooling, using the best epsilon value of 0.0001, distance weight of 1, and entropy weight of 0.001 on the Cora dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DE) 10.63 73.82 51.30 57.10
550 DMoN–DPR (DE) 10.10 72.15 41.06 44.04
243 DMoN–DPR (DE) 9.68 73.25 47.62 55.30
16 DMoN–DPR (DE) 10.61 72.13 46.23 49.09
716 DMoN–DPR (DE) 10.63 70.90 39.37 43.28
383 DMoN–DPR (DE) 9.53 74.04 47.93 54.43
277 DMoN–DPR (DE) 10.89 72.25 39.14 42.53
274 DMoN–DPR (DE) 9.49 74.39 48.31 52.84
188 DMoN–DPR (DE) 10.52 73.68 48.16 54.61
796 DMoN–DPR (DE) 11.18 72.21 42.71 45.56
Mean 10.33 72.88 45.18 49.88
Variance 0.00 0.01 0.18 0.32
Table 11: Results obtained by using DMoN–DPR pooling, using the best variance weight of 1, and entropy weight of 0.001 on the Cora dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (VE) 10.86 72.34 51.43 58.35
550 DMoN–DPR (VE) 11.10 70.30 40.10 43.20
243 DMoN–DPR (VE) 11.71 71.05 47.46 54.42
16 DMoN–DPR (VE) 11.06 70.19 45.89 48.85
716 DMoN–DPR (VE) 11.01 69.83 40.77 41.82
383 DMoN–DPR (VE) 12.14 69.28 42.65 50.92
277 DMoN–DPR (VE) 11.46 70.98 37.40 37.53
274 DMoN–DPR (VE) 11.14 72.47 46.72 48.43
188 DMoN–DPR (VE) 13.11 70.78 45.36 52.19
796 DMoN–DPR (VE) 12.85 70.48 43.74 45.16
Mean 11.64 70.77 44.15 48.09
Variance 0.01 0.01 0.17 0.39
Table 12: Results obtained by using DMoN–DPR pooling, using the best epsilon value of 0.0001, variance weight of 0.1, and entropy weight of 0.001 on the Cora dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DVE) 10.06 73.55 51.62 58.48
550 DMoN–DPR (DVE) 10.16 72.04 41.47 44.30
243 DMoN–DPR (DVE) 9.87 72.96 47.86 55.57
16 DMoN–DPR (DVE) 10.76 71.64 46.63 49.25
716 DMoN–DPR (DVE) 10.72 70.81 39.64 43.32
383 DMoN–DPR (DVE) 9.93 73.75 48.14 54.86
277 DMoN–DPR (DVE) 10.89 71.99 39.10 42.50
274 DMoN–DPR (DVE) 9.83 73.94 47.81 52.85
188 DMoN–DPR (DVE) 10.42 73.73 48.53 54.50
796 DMoN–DPR (DVE) 11.41 72.05 43.51 44.96
Mean 10.41 72.65 45.43 50.06
Variance 0.00 0.01 0.18 0.35
B.2 CiteSeer
The following tables list the performance of DMoN and different DMoN–DPR configurations across different seeds on the CiteSeer dataset.
Table 13: Results obtained by using DMoN pooling on the CiteSeer dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN 6.09 74.10 22.56 26.60
550 DMoN 4.31 75.22 25.92 41.43
243 DMoN 4.90 78.26 34.20 46.62
16 DMoN 5.87 75.48 28.89 38.28
716 DMoN 3.60 76.57 38.40 52.01
383 DMoN 4.94 75.34 26.50 39.64
277 DMoN 6.50 75.67 28.87 37.96
274 DMoN 4.35 74.93 32.36 47.53
188 DMoN 4.50 76.41 34.86 46.98
796 DMoN 4.13 75.24 30.07 44.67
Mean 4.92 75.72 30.26 42.17
Variance 0.01 0.01 0.23 0.51
Table 14: Results obtained by using DMoN–DPR pooling using the best epsilon value of 0.0001, and distance weight of 1 on the CiteSeer dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (D) 6.30 74.07 22.96 27.06
550 DMoN–DPR (D) 4.11 75.45 26.52 42.31
243 DMoN–DPR (D) 3.98 76.17 36.63 51.64
16 DMoN–DPR (D) 5.36 76.07 29.32 42.93
716 DMoN–DPR (D) 4.17 76.07 39.05 51.65
383 DMoN–DPR (D) 5.18 75.60 26.83 37.54
277 DMoN–DPR (D) 5.16 75.30 29.25 40.17
274 DMoN–DPR (D) 4.28 75.19 32.02 46.99
188 DMoN–DPR (D) 4.70 76.60 34.34 45.97
796 DMoN–DPR (D) 4.24 75.92 29.15 41.29
Mean 4.75 75.64 30.61 42.76
Variance 0.01 0.00 0.24 0.52
Table 15: Results obtained by using DMoN–DPR pooling using the best variance weight of 0.1 on the CiteSeer dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (V) 6.33 73.98 23.33 26.50
550 DMoN–DPR (V) 4.17 75.20 26.64 42.18
243 DMoN–DPR (V) 4.99 78.15 34.08 46.34
16 DMoN–DPR (V) 6.02 75.21 29.38 38.80
716 DMoN–DPR (V) 3.71 76.45 38.70 52.38
383 DMoN–DPR (V) 5.16 74.60 26.80 39.65
277 DMoN–DPR (V) 5.84 75.26 29.13 39.42
274 DMoN–DPR (V) 4.42 74.79 31.87 46.92
188 DMoN–DPR (V) 4.33 76.39 34.76 47.14
796 DMoN–DPR (V) 4.33 75.87 29.92 43.30
Mean 4.93 75.59 30.46 42.26
Variance 0.01 0.01 0.20 0.49
Table 16: Results obtained by using DMoN–DPR pooling using the best entropy weight of 0.01 on the CiteSeer dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (E) 6.59 74.00 22.17 27.07
550 DMoN–DPR (E) 4.42 75.06 26.32 41.61
243 DMoN–DPR (E) 5.14 77.95 34.48 46.79
16 DMoN–DPR (E) 6.15 75.09 29.46 38.54
716 DMoN–DPR (E) 3.67 76.36 39.12 52.80
383 DMoN–DPR (E) 5.32 74.52 27.05 39.70
277 DMoN–DPR (E) 5.89 75.27 29.14 39.34
274 DMoN–DPR (E) 4.35 74.85 32.18 47.28
188 DMoN–DPR (E) 4.55 76.16 34.69 47.23
796 DMoN–DPR (E) 4.11 75.02 30.58 45.06
Mean 5.02 75.43 30.52 42.54
Variance 0.01 0.01 0.24 0.50
Table 17: Results obtained by using DMoN–DPR pooling using the best epsilon value of 0.0001, distance weight of 1, and variance weight of 0.1 on the CiteSeer dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DV) 6.33 73.81 23.64 27.24
550 DMoN–DPR (DV) 4.09 75.48 26.65 42.08
243 DMoN–DPR (DV) 3.89 76.17 35.67 51.00
16 DMoN–DPR (DV) 5.71 75.73 29.33 42.91
716 DMoN–DPR (DV) 4.11 76.03 38.91 51.73
383 DMoN–DPR (DV) 5.16 75.31 27.06 38.22
277 DMoN–DPR (DV) 5.01 75.32 29.81 40.68
274 DMoN–DPR (DV) 4.35 74.92 32.29 47.22
188 DMoN–DPR (DV) 4.50 76.36 34.12 46.47
796 DMoN–DPR (DV) 4.44 75.93 28.46 40.94
Mean 4.76 75.51 30.59 42.85
Variance 0.01 0.01 0.21 0.50
Table 18: Results obtained by using DMoN–DPR pooling using the best epsilon value of 0.0001, distance weight of 1, and entropy weight of 0.01 on the CiteSeer dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DE) 6.48 73.71 23.56 27.55
550 DMoN–DPR (DE) 4.31 75.24 26.29 41.96
243 DMoN–DPR (DE) 3.95 76.11 36.10 51.46
16 DMoN–DPR (DE) 5.76 75.68 29.78 43.37
716 DMoN–DPR (DE) 4.15 76.00 39.03 51.88
383 DMoN–DPR (DE) 5.43 75.09 27.51 38.72
277 DMoN–DPR (DE) 5.10 75.25 29.91 40.77
274 DMoN–DPR (DE) 4.42 74.89 31.85 46.74
188 DMoN–DPR (DE) 4.75 76.28 34.05 46.33
796 DMoN–DPR (DE) 4.26 75.21 29.63 43.94
Mean 4.86 75.35 30.77 43.27
Variance 0.01 0.01 0.22 0.49
Table 19: Results obtained by using DMoN–DPR pooling using the best variance weight of 0.1, and best entropy weight of 0.01 on the CiteSeer dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (VE) 6.50 73.55 22.91 27.80
550 DMoN–DPR (VE) 4.59 74.88 26.41 41.35
243 DMoN–DPR (VE) 5.25 77.84 34.60 47.05
16 DMoN–DPR (VE) 5.95 76.05 31.65 40.71
716 DMoN–DPR (VE) 3.84 76.18 39.28 52.93
383 DMoN–DPR (VE) 5.40 74.12 26.44 39.40
277 DMoN–DPR (VE) 5.91 75.11 29.14 39.27
274 DMoN–DPR (VE) 4.55 74.52 32.02 47.12
188 DMoN–DPR (VE) 4.70 75.85 34.92 47.37
796 DMoN–DPR (VE) 4.24 74.50 31.26 45.61
Mean 5.09 75.26 30.86 42.86
Variance 0.01 0.02 0.23 0.47
Table 20: Results obtained by using DMoN–DPR pooling using the best epsilon value of 0.0001, distance weight of 1, variance weight of 0.1, and entropy weight of 0.01 on the CiteSeer dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DVE) 6.39 73.36 23.60 28.45
550 DMoN–DPR (DVE) 4.35 75.08 26.35 41.77
243 DMoN–DPR (DVE) 3.93 75.80 36.58 52.14
16 DMoN–DPR (DVE) 5.38 76.03 28.50 41.56
716 DMoN–DPR (DVE) 4.22 75.83 38.91 51.92
383 DMoN–DPR (DVE) 5.34 74.92 28.02 39.07
277 DMoN–DPR (DVE) 5.34 75.06 29.87 40.58
274 DMoN–DPR (DVE) 4.46 74.65 31.76 46.98
188 DMoN–DPR (DVE) 4.68 75.90 34.46 47.48
796 DMoN–DPR (DVE) 4.44 74.79 29.69 44.27
Mean 4.85 75.14 30.77 43.42
Variance 0.01 0.01 0.22 0.48
B.3 PubMed
The following tables list the performance of DMoN and different DMoN–DPR configurations across different seeds on the PubMed dataset.
Table 21: Results obtained by using DMoN pooling on the PubMed dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN 10.64 55.82 21.92 39.57
550 DMoN 7.91 58.30 26.08 48.88
243 DMoN 7.87 57.66 21.90 44.67
16 DMoN 8.02 57.40 21.90 44.83
716 DMoN 8.02 57.35 21.98 45.16
383 DMoN 12.56 53.65 26.03 49.14
277 DMoN 7.78 57.61 21.73 44.90
274 DMoN 7.94 57.44 21.59 44.78
188 DMoN 7.90 57.50 21.86 45.03
796 DMoN 7.93 57.53 21.76 44.82
Mean 8.66 57.03 22.68 45.18
Variance 0.03 0.02 0.03 0.07
Table 22: Results obtained by using DMoN–DPR pooling using the best epsilon value of 0.001 and distance weight of 1 on the PubMed dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (D) 10.63 55.83 21.92 39.59
550 DMoN–DPR (D) 7.94 58.27 26.16 48.94
243 DMoN–DPR (D) 7.88 57.65 21.93 44.67
16 DMoN–DPR (D) 8.00 57.41 21.80 44.75
716 DMoN–DPR (D) 8.07 57.30 21.97 45.11
383 DMoN–DPR (D) 12.53 53.69 25.84 49.03
277 DMoN–DPR (D) 7.79 57.60 21.73 44.91
274 DMoN–DPR (D) 7.96 57.43 21.76 44.94
188 DMoN–DPR (D) 7.91 57.49 21.87 45.05
796 DMoN–DPR (D) 7.92 57.54 21.84 44.89
Mean 8.66 57.02 22.68 45.19
Variance 0.03 0.02 0.03 0.07
Table 23: Results obtained by using DMoN–DPR pooling using the best variance weight of 0.001 on the PubMed dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (V) 10.64 55.82 21.91 39.57
550 DMoN–DPR (V) 7.92 58.29 26.08 48.88
243 DMoN–DPR (V) 7.88 57.65 21.91 44.68
16 DMoN–DPR (V) 8.01 57.40 21.92 44.86
716 DMoN–DPR (V) 8.02 57.35 21.97 45.16
383 DMoN–DPR (V) 12.56 53.64 26.03 49.14
277 DMoN–DPR (V) 7.78 57.61 21.75 44.89
274 DMoN–DPR (V) 7.94 57.44 21.59 44.78
188 DMoN–DPR (V) 7.91 57.49 21.86 45.04
796 DMoN–DPR (V) 7.92 57.54 21.78 44.83
Mean 8.66 57.02 22.68 45.18
Variance 0.03 0.02 0.03 0.07
Table 24: Results obtained by using DMoN–DPR pooling using the best entropy weight of 0.001 on the PubMed dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (E) 10.64 55.82 21.91 39.59
550 DMoN–DPR (E) 7.92 58.29 26.08 48.88
243 DMoN–DPR (E) 7.88 57.65 21.90 44.67
16 DMoN–DPR (E) 8.01 57.40 21.90 44.83
716 DMoN–DPR (E) 8.03 57.34 21.96 45.13
383 DMoN–DPR (E) 12.55 53.65 26.04 49.15
277 DMoN–DPR (E) 7.78 57.61 21.75 44.91
274 DMoN–DPR (E) 7.96 57.42 21.63 44.80
188 DMoN–DPR (E) 7.93 57.47 21.88 45.06
796 DMoN–DPR (E) 7.95 57.52 21.79 44.85
Mean 8.67 57.02 22.68 45.19
Variance 0.03 0.02 0.03 0.07
Table 25: Results obtained by using DMoN–DPR pooling using the best epsilon value of 0.001, distance weight of 1, and variance weight of 0.001 on the PubMed dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DV) 10.63 55.84 21.89 39.58
550 DMoN–DPR (DV) 7.93 58.27 26.15 48.92
243 DMoN–DPR (DV) 7.89 57.64 21.94 44.67
16 DMoN–DPR (DV) 7.97 57.43 21.83 44.78
716 DMoN–DPR (DV) 8.06 57.31 22.00 45.11
383 DMoN–DPR (DV) 12.52 53.69 25.81 49.01
277 DMoN–DPR (DV) 7.79 57.59 21.72 44.91
274 DMoN–DPR (DV) 7.93 57.46 21.73 44.92
188 DMoN–DPR (DV) 7.91 57.48 21.83 45.03
796 DMoN–DPR (DV) 7.93 57.53 21.86 44.93
Mean 8.66 57.02 22.68 45.19
Variance 0.03 0.02 0.03 0.07
Table 26: Results obtained by using DMoN–DPR pooling using the best epsilon value of 0.001, distance weight of 1, and entropy weight of 0.001 on the PubMed dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DE) 10.60 55.87 21.90 39.66
550 DMoN–DPR (DE) 7.94 58.27 26.15 48.91
243 DMoN–DPR (DE) 7.89 57.64 21.93 44.67
16 DMoN–DPR (DE) 7.97 57.43 21.83 44.78
716 DMoN–DPR (DE) 8.08 57.29 22.01 45.14
383 DMoN–DPR (DE) 12.54 53.68 25.80 49.00
277 DMoN–DPR (DE) 7.79 57.59 21.71 44.90
274 DMoN–DPR (DE) 7.97 57.42 21.75 44.94
188 DMoN–DPR (DE) 7.93 57.47 21.86 45.01
796 DMoN–DPR (DE) 7.92 57.54 21.78 44.87
Mean 8.66 57.02 22.67 45.19
Variance 0.03 0.02 0.03 0.07
Table 27: Results obtained by using DMoN–DPR pooling using the best variance weight of 0.001, and entropy weight of 0.001 on the PubMed dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (VE) 10.64 55.83 21.89 39.57
550 DMoN–DPR (VE) 7.92 58.29 26.09 48.88
243 DMoN–DPR (VE) 7.87 57.65 21.89 44.67
16 DMoN–DPR (VE) 8.01 57.40 21.91 44.85
716 DMoN–DPR (VE) 8.03 57.33 21.96 45.13
383 DMoN–DPR (VE) 12.57 53.63 26.02 49.14
277 DMoN–DPR (VE) 7.78 57.60 21.74 44.91
274 DMoN–DPR (VE) 7.96 57.42 21.61 44.81
188 DMoN–DPR (VE) 7.92 57.47 21.90 45.03
796 DMoN–DPR (VE) 7.94 57.52 21.77 44.84
Mean 8.66 57.01 22.68 45.18
Variance 0.03 0.02 0.03 0.07
Table 28: Results obtained by using DMoN–DPR pooling using the best epsilon value of 0.001, distance weight of 1, variance weight of 0.001, and entropy weight of 0.001 on the PubMed dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DVE) 10.59 55.87 21.90 39.65
550 DMoN–DPR (DVE) 7.93 58.27 26.14 48.91
243 DMoN–DPR (DVE) 7.89 57.64 21.93 44.67
16 DMoN–DPR (DVE) 7.97 57.42 21.84 44.78
716 DMoN–DPR (DVE) 8.08 57.29 21.91 45.06
383 DMoN–DPR (DVE) 12.51 53.69 25.79 48.98
277 DMoN–DPR (DVE) 7.79 57.58 21.70 44.88
274 DMoN–DPR (DVE) 7.94 57.45 21.74 44.94
188 DMoN–DPR (DVE) 7.94 57.45 21.89 45.04
796 DMoN–DPR (DVE) 7.92 57.54 21.81 44.89
Mean 8.66 57.02 22.67 45.18
Variance 0.03 0.02 0.03 0.07
B.4 Coauthor CS
The following tables list the performance of DMoN and different DMoN–DPR configurations across different seeds on the Coauthor CS dataset.
Table 29: Results obtained by using DMoN pooling on the Coauthor CS dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN 19.07 72.23 68.48 58.27
550 DMoN 19.14 71.64 66.78 55.19
243 DMoN 18.11 72.88 70.78 60.88
16 DMoN 18.28 73.04 71.13 62.23
716 DMoN 18.41 72.63 69.20 59.44
383 DMoN 18.45 72.43 69.96 59.60
277 DMoN 19.53 72.25 67.51 54.79
274 DMoN 20.20 71.69 69.38 58.09
188 DMoN 17.89 73.05 71.52 60.61
796 DMoN 18.89 72.50 69.77 59.18
Mean 18.80 72.43 69.45 58.83
Variance 0.01 0.00 0.02 0.06
Table 30: Results obtained by using DMoN–DPR pooling using the best epsilon value of 1, and distance weight of 1 on the Coauthor CS dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR(D) 18.88 72.53 72.11 62.71
550 DMoN–DPR(D) 20.80 71.19 68.18 57.76
243 DMoN–DPR(D) 19.36 71.83 71.19 62.87
16 DMoN–DPR(D) 18.99 72.42 71.12 63.30
716 DMoN–DPR(D) 17.85 73.08 70.14 56.54
383 DMoN–DPR(D) 19.24 71.60 69.23 60.49
277 DMoN–DPR(D) 18.20 73.02 72.35 64.04
274 DMoN–DPR(D) 18.86 72.11 70.97 61.56
188 DMoN–DPR(D) 19.85 71.18 68.61 59.59
796 DMoN–DPR(D) 18.02 72.72 71.70 62.08
Mean 19.01 72.17 70.56 61.09
Variance 0.01 0.00 0.02 0.06
Table 31: Results obtained by using DMoN–DPR pooling using the best variance weight of 0.1 on the Coauthor CS dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (V) 18.87 72.35 68.79 58.77
550 DMoN–DPR (V) 19.03 71.68 67.07 56.15
243 DMoN–DPR (V) 18.27 72.74 71.13 61.15
16 DMoN–DPR (V) 18.28 72.94 71.38 62.48
716 DMoN–DPR (V) 19.77 71.86 68.11 57.05
383 DMoN–DPR (V) 18.53 72.22 70.15 60.21
277 DMoN–DPR (V) 19.78 72.14 67.97 55.30
274 DMoN–DPR (V) 19.57 72.49 70.45 59.36
188 DMoN–DPR (V) 18.24 72.89 71.02 59.87
796 DMoN–DPR (V) 18.52 72.76 71.57 60.67
Mean 18.89 72.41 69.76 59.10
Variance 0.00 0.00 0.03 0.05
Table 32: Results obtained by using DMoN–DPR pooling using the best entropy weight of 0.1 on the Coauthor CS dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (E) 18.95 71.61 72.94 59.97
550 DMoN–DPR (E) 19.87 71.23 70.05 56.24
243 DMoN–DPR (E) 20.13 70.45 71.39 60.47
16 DMoN–DPR (E) 19.44 71.06 72.47 64.02
716 DMoN–DPR (E) 19.11 71.28 73.55 64.42
383 DMoN–DPR (E) 20.09 70.20 67.61 54.84
277 DMoN–DPR (E) 21.27 69.71 68.97 56.84
274 DMoN–DPR (E) 20.01 71.45 73.13 62.46
188 DMoN–DPR (E) 18.70 71.68 72.18 63.95
796 DMoN–DPR (E) 19.74 71.04 71.06 60.77
Mean 19.73 70.97 71.34 60.40
Variance 0.01 0.00 0.04 0.12
Table 33: Results obtained by using DMoN–DPR pooling using the best epsilon value of 1, distance weight of 1, and variance weight of 0.1 on the Coauthor CS dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DV) 18.76 72.48 72.77 63.41
550 DMoN–DPR (DV) 19.71 71.95 70.36 61.52
243 DMoN–DPR (DV) 19.10 72.09 71.75 63.65
16 DMoN–DPR (DV) 18.57 72.77 73.03 63.98
716 DMoN–DPR (DV) 19.02 72.54 70.44 58.08
383 DMoN–DPR (DV) 18.57 72.33 72.29 63.93
277 DMoN–DPR (DV) 17.94 73.14 72.46 63.13
274 DMoN–DPR (DV) 19.41 71.73 70.41 60.45
188 DMoN–DPR (DV) 20.33 71.18 69.03 58.99
796 DMoN–DPR (DV) 18.06 72.38 70.79 60.99
Mean 18.95 72.26 71.33 61.81
Variance 0.01 0.00 0.02 0.05
Table 34: Results obtained by using DMoN–DPR pooling using the best epsilon value of 1, distance weight of 1, and entropy weight of 0.1 on the Coauthor CS dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DE) 19.82 70.82 71.87 64.34
550 DMoN–DPR (DE) 19.72 70.51 70.06 61.80
243 DMoN–DPR (DE) 19.74 70.31 70.33 61.84
16 DMoN–DPR (DE) 19.83 71.73 74.31 63.08
716 DMoN–DPR (DE) 18.97 71.30 71.50 62.63
383 DMoN–DPR (DE) 19.97 70.74 72.47 65.26
277 DMoN–DPR (DE) 19.50 71.15 72.02 60.54
274 DMoN–DPR (DE) 19.42 69.84 68.51 59.93
188 DMoN–DPR (DE) 21.62 69.25 67.61 56.77
796 DMoN–DPR (DE) 20.19 69.93 69.69 61.41
Mean 19.88 70.56 70.84 61.76
Variance 0.00 0.01 0.04 0.06
Table 35: Results obtained by using DMoN–DPR pooling using the best variance weight of 0.1, and entropy weight of 0.1 on the Coauthor CS dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (VE) 19.08 71.49 73.11 60.17
550 DMoN–DPR (VE) 19.95 71.16 70.08 56.20
243 DMoN–DPR (VE) 20.21 70.35 71.26 60.33
16 DMoN–DPR (VE) 19.62 70.89 72.13 63.50
716 DMoN–DPR (VE) 19.16 71.25 73.42 64.30
383 DMoN–DPR (VE) 20.23 70.06 67.62 54.87
277 DMoN–DPR (VE) 20.14 70.81 70.43 59.06
274 DMoN–DPR (VE) 20.04 71.40 73.23 62.31
188 DMoN–DPR (VE) 18.72 71.67 72.17 63.93
796 DMoN–DPR (VE) 19.80 70.97 71.09 60.92
Mean 19.70 71.01 71.45 60.56
Variance 0.00 0.00 0.03 0.10
Table 36: Results obtained by using DMoN–DPR pooling using the best epsilon value of 1, distance weight of 1, variance weight of 0.1, and entropy weight of 0.1 on the Coauthor CS dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DVE) 19.32 71.13 72.76 65.12
550 DMoN–DPR (DVE) 19.75 70.50 70.00 61.78
243 DMoN–DPR (DVE) 19.97 70.19 70.40 61.86
16 DMoN–DPR (DVE) 18.87 72.23 74.23 65.17
716 DMoN–DPR (DVE) 19.03 71.26 71.56 62.76
383 DMoN–DPR (DVE) 20.06 70.65 72.47 65.18
277 DMoN–DPR (DVE) 19.06 72.07 73.97 64.36
274 DMoN–DPR (DVE) 19.41 69.81 68.57 60.03
188 DMoN–DPR (DVE) 21.66 69.20 67.65 56.86
796 DMoN–DPR (DVE) 20.30 69.78 69.21 60.98
Mean 19.74 70.68 71.08 62.41
Variance 0.01 0.01 0.05 0.07
B.5 Coauthor Physics
The following tables list the performance of DMoN and different DMoN–DPR configurations across different seeds on the Coauthor Physics dataset.
Table 37: Results obtained by using DMoN pooling on the Coauthor Physics dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN 12.61 60.14 44.16 46.45
550 DMoN 14.36 63.58 51.07 45.68
243 DMoN 13.84 63.82 56.68 49.40
16 DMoN 13.81 63.78 57.00 49.64
716 DMoN 13.91 63.70 55.59 49.16
383 DMoN 13.84 63.08 48.79 45.27
277 DMoN 14.06 63.70 46.72 39.63
274 DMoN 13.84 63.93 53.70 46.41
188 DMoN 13.94 63.70 55.66 49.31
796 DMoN 13.89 64.04 54.63 47.03
Mean 13.81 63.35 52.40 46.80
Variance 0.00 0.01 0.20 0.09
Table 38: Results obtained by using DMoN–DPR pooling using the best epsilon value of 10, and distance weight of 1 on the Coauthor Physics dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (D) 16.08 60.64 49.75 44.05
550 DMoN–DPR (D) 15.42 56.02 50.59 52.21
243 DMoN–DPR (D) 15.78 58.21 58.58 53.21
16 DMoN–DPR (D) 15.91 54.69 53.21 54.59
716 DMoN–DPR (D) 15.85 57.35 50.14 49.28
383 DMoN–DPR (D) 18.82 55.05 51.96 48.46
277 DMoN–DPR (D) 19.89 54.78 50.10 47.08
274 DMoN–DPR (D) 15.03 56.06 55.73 54.53
188 DMoN–DPR (D) 13.71 59.75 61.71 54.99
796 DMoN–DPR (D) 18.02 56.74 58.17 51.22
Mean 16.45 56.93 53.99 50.96
Variance 0.04 0.04 0.18 0.13
Table 39: Results obtained by using DMoN–DPR pooling using the best variance weight of 1 on the Coauthor Physics dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (V) 12.00 58.02 45.50 45.54
550 DMoN–DPR (V) 12.44 57.44 53.99 54.11
243 DMoN–DPR (V) 13.82 61.60 59.00 51.98
16 DMoN–DPR (V) 13.64 58.69 58.90 56.75
716 DMoN–DPR (V) 14.06 61.38 56.03 51.89
383 DMoN–DPR (V) 14.42 63.20 52.72 45.15
277 DMoN–DPR (V) 12.96 57.38 52.20 53.30
274 DMoN–DPR (V) 12.88 57.23 54.09 56.20
188 DMoN–DPR (V) 13.95 57.14 52.27 54.06
796 DMoN–DPR (V) 14.29 61.71 55.55 49.94
Mean 13.45 59.38 54.03 51.89
Variance 0.01 0.05 0.15 0.16
Table 40: Results obtained by using DMoN–DPR pooling using the best entropy weight of 0.1 on the Coauthor Physics dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (E) 12.88 58.41 44.16 45.67
550 DMoN–DPR (E) 12.07 57.62 54.04 53.83
243 DMoN–DPR (E) 13.79 62.02 58.66 51.38
16 DMoN–DPR (E) 13.77 60.22 58.66 53.25
716 DMoN–DPR (E) 14.41 61.97 56.14 50.58
383 DMoN–DPR (E) 13.88 58.47 53.74 53.85
277 DMoN–DPR (E) 12.85 57.66 52.26 52.98
274 DMoN–DPR (E) 15.10 59.77 51.39 45.72
188 DMoN–DPR (E) 14.12 62.20 56.74 50.19
796 DMoN–DPR (E) 15.15 62.10 45.87 43.09
Mean 13.80 60.04 53.17 50.05
Variance 0.01 0.04 0.25 0.15
Table 41: Results obtained by using DMoN–DPR pooling using the best epsilon value of 10, distance weight of 1, and variance weight 1 on the Coauthor Physics dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DV) 14.08 60.60 47.53 46.00
550 DMoN–DPR (DV) 14.07 55.94 54.18 57.79
243 DMoN–DPR (DV) 13.03 58.15 60.48 59.90
16 DMoN–DPR (DV) 12.03 55.51 54.29 57.77
716 DMoN–DPR (DV) 12.54 56.50 52.40 57.28
383 DMoN–DPR (DV) 13.82 56.10 52.59 57.34
277 DMoN–DPR (DV) 13.96 56.84 53.28 57.82
274 DMoN–DPR (DV) 12.40 56.41 52.47 57.72
188 DMoN–DPR (DV) 14.36 56.80 60.08 57.68
796 DMoN–DPR (DV) 14.34 55.61 53.64 54.26
Mean 13.46 56.85 54.09 56.36
Variance 0.01 0.02 0.14 0.15
Table 42: Results obtained by using DMoN–DPR pooling using the best epsilon value of 10, distance weight of 1, and entropy weight of 0.1 on the Coauthor Physics dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DE) 15.49 55.62 48.15 53.66
550 DMoN–DPR (DE) 15.31 55.21 52.10 56.57
243 DMoN–DPR (DE) 13.74 56.71 59.84 58.07
16 DMoN–DPR (DE) 12.99 55.20 53.27 56.69
716 DMoN–DPR (DE) 14.00 56.38 59.34 58.29
383 DMoN–DPR (DE) 13.94 56.83 55.09 56.84
277 DMoN–DPR (DE) 14.97 54.26 54.14 60.85
274 DMoN–DPR (DE) 12.57 55.85 57.05 59.56
188 DMoN–DPR (DE) 12.99 56.54 58.60 58.86
796 DMoN–DPR (DE) 16.16 58.41 49.49 49.99
Mean 14.22 56.10 54.71 56.94
Variance 0.01 0.01 0.16 0.10
Table 43: Results obtained by using DMoN–DPR pooling using the best variance weight of 1, and entropy weight 0f 0.1 on the Coauthor Physics dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (VE) 8.50 36.32 14.32 44.25
550 DMoN–DPR (VE) 13.12 57.28 52.98 56.87
243 DMoN–DPR (VE) 14.64 59.71 55.12 51.50
16 DMoN–DPR (VE) 13.92 58.45 62.37 58.80
716 DMoN–DPR (VE) 13.75 57.27 52.06 51.47
383 DMoN–DPR (VE) 14.04 57.77 56.72 58.23
277 DMoN–DPR (VE) 13.04 57.90 54.22 57.55
274 DMoN–DPR (VE) 12.10 56.56 51.63 56.60
188 DMoN–DPR (VE) 10.40 55.47 55.62 62.42
796 DMoN–DPR (VE) 15.01 60.44 46.37 47.48
Mean 12.85 55.72 50.14 54.52
Variance 0.04 0.49 1.75 0.32
Table 44: Results obtained by using DMoN–DPR pooling using the best epsilon value of 10, distance weight of 1, variance weight of 1, and entropy weight of 0.1 on the Coauthor Physics dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DVE) 15.43 55.58 47.68 53.96
550 DMoN–DPR (DVE) 14.27 55.01 53.99 58.87
243 DMoN–DPR (DVE) 12.76 56.10 53.71 58.34
16 DMoN–DPR (DVE) 14.62 53.36 49.11 57.45
716 DMoN–DPR (DVE) 13.06 54.46 57.73 62.18
383 DMoN–DPR (DVE) 12.92 55.12 58.39 60.97
277 DMoN–DPR (DVE) 14.04 55.08 51.68 58.26
274 DMoN–DPR (DVE) 12.59 55.68 58.34 60.52
188 DMoN–DPR (DVE) 11.41 57.40 63.84 62.83
796 DMoN–DPR (DVE) 13.51 54.94 48.67 55.26
Mean 13.46 55.27 54.31 58.86
Variance 0.01 0.01 0.27 0.08
Appendix C Code Implementation
The code was implemented by extending the DMoN implementation in PyTorch Geometric (Fey and Lenssen, 2019), and was trained and evaluated using the evaluation protocols found in the official DMoN repository (Tsitsulin et al., 2023).